What is the x- coordinate of the solution to the system of linear equations 3X+6Y=4 AND X-3Y=1,
?

Answer by ewatrrr(24785)


Hi,
finding the x- coordinate of the solution to following system
3X+6Y=4 OR y = (-3/6)x + 4/6
X-3Y=1 OR y = (1/3)x - 1/3
when these lines intersect, the y-value will be the same for each.
(-3/6)x + 4/6 = 1/3x -1/3 Multiplying each term by 6
-3x + 4 = 2x - 2
6 = 5x
6/5 = x
CHECKING our Answer: Note: y = 1/3*6/5 -1/3 = 1/15
y = (-3/6)x + 4/6
1/15 = (-3/6)*6/5 + 4/6 = -18/30 + 20/30 = 2/30 = 1/15